### v1

# Intertwined order of generalized global symmetries

Abstract
We investigate the interplay of generalized global symmetries in 2+1 dimensions by introducing a lattice model that couples a $\mathbb{Z}_{N}$ clock model to a $\mathbb{Z}_{N}$ gauge theory via a topological interaction. This coupling binds the charges of one symmetry to the disorder operators of the other, and when these composite objects condense, they give rise to emergent generalized symmetries with mixed ’t Hooft anomalies. These anomalies result in phases with ordinary symmetry breaking, topological order, and symmetry-protected topological (SPT) order, where the different types of order are not independent but intimately related. We further explore the gapped boundary states of these exotic phases and develop theories for phase transitions between them. Additionally, we extend our lattice model to incorporate a non-invertible global symmetry, which can be spontaneously broken, leading to domain walls with non-trivial fusion rules.
1 Introduction
Symmetry plays an fundamental role in physics and particularly in condensed matter physics. In the conventional Landau theory of phase transitions, phases of matter are classified by the global symmetries they spontaneously break. On the other hand, symmetry-protected topological phases (SPTs) are characterized by their nontrivial response to probes of global symmetries or by the ’t Hooft anomalies of their boundary states [1, 2, 3, 4, 5, 6].
Recent developments have expanded the concept of global symmetry in several ways, thus widening our perspective on how to characterize phases of matter [7, 8, 9, 10]. While an ordinary global symmetry acts on local operators, a $q$-form global symmetry acts on operators supported on $q$-dimensional objects. Using this terminology, an ordinary global symmetry is a zero-form symmetry. We will be especially interested in one-form global symmetries, which act on operators defined on noncontractible loops, since these symmetries can be used to characterize topologically ordered phases [11, 12]. Indeed, the braiding of two Abelian anyons in (2+1)d results in a phase factor, which can be interpreted as resulting from the action of a one-form symmetry by the worldline of one anyon on another. A phase with Abelian topological order may then be thought of as “spontaneously” breaking a discrete one-form global symmetry in the sense that the symmetry is respected by the low energy effective action (or Hamiltonian) but not by a given ground state [^1] For ordinary zero-form global symmetries, the precise definition of spontaneous symmetry breaking is that upon coupling to a local external symmetry-breaking field $h(x)$, the order parameter remains nonzero if we take the thermodynamic limit and then $h\to 0$. A analogue of this criterion does not exist for higher-form symmetries because the observables are non-local. However, a spontaneously broken one-form global symmetry can be defined as the condition that the correlators of Wilson operators on noncontractible loops obey clustering. A well-known example is the Polyakov loop used to diagnose the thermal phase transition from a confined to a deconfined phase of a pure gauge theory [13].. We emphasize that this one-form symmetry is an emergent infrared (IR) symmetry in the usual condensed matter setting, but once it emerges, it is extremely robust [14].
A distinct generalization of symmetry that has garnered much attention recently is non-invertible symmetry [15, 16]. Conventional symmetry operators are unitary (or antiunitary) and thus have inverses. Recently, it has been argued that topological operators that do not necessarily have inverses should be viewed as symmetry operators [17, 18]. While ordinary global symmetries are characterized mathematically by groups, symmetry operators for non-invertible symmetries can have multiple fusion channels. A flurry of recent work has introduced many examples of field theories [19, 20, 21, 22] and lattice models [23, 24, 25] with non-invertible symmetries, demonstrating that these symmetries can lead to novel SPTs [26, 27] and intricate phase diagrams with exotic symmetry-breaking patterns [28, 29, 30].
Experience with ordinary global symmetries suggests that the interplay of different symmetries can give rise to rich phase diagrams with a variety of interesting phases. In particular, the concept of intertwined order has been invoked to understand general principles at play in the complex phase diagrams of correlated electronic systems [31]. The idea is that the different orders in a complex phase diagram can arise from an underlying microscopic primary state. A paradigmatic example of a possible parent state is a pair density wave superconductor [32], which spontaneously breaks $U(1)$ particle number conservation and translation symmetry, which hosts several ordered states.
We therefore should expect that a rich set of physics to arise from models with an interplay of generalized global symmetries. In this work, we develop such models with intertwined generalized symmetries. One way in which generalized symmetries can be intertwined is that they can form a higher-group symmetry [33, 34], but we take a different approach here. With some inspiration from a previous work of the authors [35], our starting point is a 3d Euclidean lattice model consisting of a $\mathbb{Z}_{N}$ clock model coupled to a $\mathbb{Z}_{N}$ gauge theory (defined on dual lattices) by a topological term rather than by minimal coupling [^2] In this model the $\mathbb{Z}_{N}$ global symmetry of the clock model is not gauged and the $\mathbb{Z}_{N}$ local symmetry of the gauge theory is not Higgsed.. This topological coupling is analogous to the theta term in (3+1)d gauge theories, which gives magnetic monopoles an electric charge—a phenomenon known as the Witten effect [36]. Similarly, our (2+1)d topological interaction binds vortices of the spin model to an electric charge of the gauge theory, and the magnetic monopoles of the gauge theory become bound to the spins. Different combinations of these composite operators can condense, leading to a rich phase diagram with a variety of interesting phases. Using duality arguments, we can deduce much of this phase diagram and the physics of the phases in the model.
One class of phases that arise in this model are SPTs protected by the $\mathbb{Z}_{N}$ zero-form and $\mathbb{Z}_{N}$ one-form global symmetries. We denote a $\mathbb{Z}_{N}$ $q$-form global symmetry by $\mathbb{Z}_{N}^{(q)}$ so that the symmetry protecting the SPTs of this type is $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$. These SPT phases are characterized in the continuum by the response,
$S_{\mathrm{SPT}}[A_{\mu},B_{\mu\nu}]=\frac{iNp}{2\pi}\int A\wedge B=\frac{iNp}{4\pi}\int d^{3}x\,\varepsilon_{\mu\nu\lambda}\,A_{\mu}\,B_{\nu\lambda},$ (1.1)
where $A_{\mu}$ is a $\mathbb{Z}_{N}$ one-form background gauge field, $B_{\mu\nu}$ is a $\mathbb{Z}_{N}$ two-form background gauge field, and $p\in\mathbb{Z}$ is a parameter characterizing the SPT phase. A lattice model for an SPT of this type was first described for $N=2$ and $p=1$ on the triangular lattice in Ref. [37]. For a given $N$, an SPT with each possible value of $p$ is realized in a phase of the lattice model. Our lattice model thus serves as a microscopic model for every SPT phase with a response of the form in Eq. (1.1), thus providing a versatile platform to study these phases and their phase transitions.
Our lattice model also admits more intricate phases characterized by an effective field theory that is a gauged version of Eq. (1.1), meaning that $A_{\mu}$ and $B_{\mu\nu}$ are both promoted to fluctuating gauge fields. By the correspondence between SPTs and topologically ordered states [38], these phases are symmetry enriched topological states (SETs) that have topological order, but there is some additional symmetry breaking of the ordinary global symmetry. These phases are generated by simultaneously condensing bound states of local operators with spin charge $N$ and magnetic charge $p$ and loop operators with electric charge $N$ and vorticity $p$, which is reminiscent of the physics of oblique confining phases in (3+1)d gauge theories [39, 40, 41], so we refer to this class of (2+1)d phases as oblique phases. We demonstrate that for $L=\gcd(N,p)>1$, the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry has mixed ’t Hooft anomalies with emergent generalized symmetries, resulting in a breaking of this symmetry to a nontrivial subgroup, $H=\mathbb{Z}_{N/L}^{(0)}\times\mathbb{Z}_{N/L}^{(1)}$, giving rise to both ordinary symmetry breaking and topological order. Furthermore, the unbroken subgroup $H$ has a nontrivial response to background fields, signaling SPT order for this symmetry.
To emphasize the new features of our lattice model, let us contrast with another physical system that has both ordinary symmetry breaking and topological order. For example, fractional quantum Hall systems can certainly host broken symmetry states with nematic order that coexists with topological order [42, 43]. However, the symmetry structure of the systems we are studying here are more intricate. The physics described above for an oblique phase is representative of the interplay of the zero-form and one-form symmetries in our lattice model more generally. The patterns of symmetry breaking for the two symmetries are not independent and do not merely coexist. For any phase of our lattice model, if the $\mathbb{Z}_{N}$ zero-form symmetry is spontaneously broken to a nontrivial subgroup, the $\mathbb{Z}_{N}$ one-form symmetry is also broken to the same subgroup. The remaining unbroken zero-form and one-form symmetries additionally have mixed SPT order. These features demonstrate that the symmetries are not independent but indeed have a special interplay throughout the phase diagram because of the way in which they are coupled in the lattice model.
For systems with an open boundary, for every SPT of the form in Eq. (1.1), preserving the zero-form and one-form symmetries inevitably leads to boundary modes. The one-form symmetry cannot be spontaneously broken along the (1+1)d boundary, so any gapped boundary state must necessarily break the zero-form symmetry spontaneously. We develop the precise criteria the boundary theory must satisfy to cancel the ’t Hooft anomaly of the bulk and provide examples. We also develop two different gapped boundary states for oblique phases. One of these boundary conditions, which we refer to as an electric boundary condition, preserves the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry of the bulk. This boundary condition is consistent with the SPT order of the unbroken $H=\mathbb{Z}_{N/L}^{(0)}\times\mathbb{Z}_{N/L}^{(1)}$ subgroup and results in the spontaneous breaking of $H$ at the boundary. The other gapped boundary condition, which we call the magnetic boundary state, preserves a magnetic $\mathbb{Z}_{p}^{(0)}\times\mathbb{Z}_{p}^{(1)}$ symmetry, resulting in the spontaneous breaking of this symmetry at the boundary.
We also find possible gapless boundary states for $\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ SPTs. By analyzing possible perturbations that gap the boundary, we show that the gapless state can be interpreted as a quantum critical point (or line) that has additional emergent symmetries. In the simplest case, we consider the gapless boundary state of the SPT, Eq. (1.1), and add only perturbations that preserve the $\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry coupled to the bulk. For $L=\gcd(N,p)\geq 4$, our gapless state is a quantum critical point (critical line for $L>4$) where the global symmetry is enhanced to $U(1)^{(0)}\times U(1)^{(0)}\times\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$. This critical point or line then separates two gapped phases, each of which has a different $\mathbb{Z}_{N}^{(0)}$ symmetry that is spontaneously broken.
Finally, we consider generalizations of our Euclidean lattice model in which the parameter $\Theta$, the coefficient of the topological interaction, is promoted to a dynamical matter field associated with an additional $\mathbb{Z}_{N}^{(0)}$ global symmetry. We refer to this new matter field as a $\mathbb{Z}_{N}$ axion in analogy with the usual axion that couples to the theta term of a gauge theory in (3+1)d. The new $(\mathbb{Z}_{N}^{(0)})_{\mathrm{axion}}$ symmetry has a mixed ’t Hooft anomaly with the original $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry, which ensures that no phase of this lattice model is gapped and preserves all the symmetries. An especially interesting phase occurs when the $(\mathbb{Z}_{N}^{(0)})_{\mathrm{axion}}$ symmetry is spontaneously broken while the remaining $G$ symmetry is preserved. In this case, the domain walls for the $(\mathbb{Z}_{N}^{(0)})_{\mathrm{axion}}$ symmetry separate distinct $G$ SPT states.
If we now gauge the $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ symmetry of the lattice model, the domain walls will now separate distinct oblique states characterized by different topological orders, spontaneous symmetry breaking, and mixed SPT orders. As with the boundary states of the oblique phases, we must then necessarily decorate the domain walls with dynamical degrees of freedom, which is somewhat reminiscent of the decorated domain wall construction for SPTs [44], but here, the domain walls are decorated with a lower-dimensional state with spontaneous symmetry breaking rather than SPT order. The physical consequence of introducing these degrees of freedom is that the domain walls no longer obey group-like fusion rules. This generalization of our lattice model thus has a phase in which a non-invertible symmetry is spontaneously broken. Indeed, the mixed anomaly implies that gauging $G$ anomalously breaks the $(\mathbb{Z}_{N}^{(0)})_{\mathrm{axion}}$ symmetry, but we demonstrate that it can alternatively be viewed as a non-invertible symmetry. This notion of generating non-invertible symmetries from mixed anomalies as appeared in other contexts previously [19, 21].
This paper is organized as follows. In Section 2, we introduce our lattice model and discuss its global symmetries and phase diagram. Section 3 describes the low energy physics and effective field theory of the oblique phases. In Section 4, we discuss the intertwined SPT response of a generic oblique phase. Section 5 develops the phase transitions of the bulk phases of the lattice model in more detail. We then derive possible gapped boundary states of both the SPT state and the oblique phases in Section 6. Section 7 introduces a Hamiltonian lattice model that realizes oblique phases as its ground states. In Section 8, we develop gapless boundary states of the SPT states and present examples of boundary phase transitions into other gapped boundary states. Finally, in Section 9, we study the two $\mathbb{Z}_{N}$ axion generalizations of our lattice model, one of which has mixed anomalies of generalized symmetries in the UV, and another which has a non-invertible symmetry. In Appendix A, we review $\mathbb{Z}_{N}$ gauge fields in the continuum. In Appendix B, we find dualities of our lattice model. In Appendix C, we derive an effective Coulomb gas action for our lattice model and relate it to duality. Appendix D reviews the lattice version of (2+1)d BF theory and its counterpart in the ordered phase of the clock model. Appendix E contains some of the more technical calculations associated with the $\mathbb{Z}_{N}$ axion models.
2 Euclidean lattice model
2.1 Model
Figure 1: Depiction of the topological interaction on the lattice, Eq. (2.4). This (2+1)d analogue of the theta term represents an interaction of the field strengths, $f_{\nu\lambda}$, which are on plaquettes of the direct lattice, and spin phase variations, $\Delta_{\mu}\varphi$, which live on links of the dual lattice.
To study the interplay of zero-form and one-form global symmetries, we should use a model that couples these symmetries together. We thus consider a model of a $\mathbb{Z}_{N}$ lattice gauge theory coupled to $\mathbb{Z}_{N}$ matter on a three-dimensional cubic lattice in Euclidean spacetime [^3] See Refs. [45, 46, 47] for a review of lattice gauge theory.. However, the matter is not minimally coupled to the gauge field because this type of coupling removes the global symmetries we are interested in examining. Instead, we place the gauge fields $a_{\mu}$ on links of the lattice, matter fields $\varphi$ on sites of the dual lattice (i.e., centers of cubes on the direct lattice), and couple the matter to the gauge fields using a lattice analogue of a topological term. This choice of model was largely inspired by previous work of the authors [35] on the Cardy-Rabinovici model [40, 41, 48, 49, 50], a (3+1)d $\mathbb{Z}_{N}$ lattice gauge theory with an analogue of a theta term. Here, we consider a partition function in the Villain form [51], given by
$\begin{split}Z&=\int[da_{\mu}][d\varphi]\sum_{\left\{n,\,n_{\mu},\,s_{\mu},\,s_{\mu\nu}\right\}}\delta(\Delta_{\mu}n_{\mu})\,e^{-S[\varphi,\,a_{\mu},\,n,\,n_{\mu},\,s_{\mu},\,s_{\mu\nu}]},\\ S&=\frac{1}{4e^{2}}\sum_{P}(\Delta_{\mu}a_{\nu}-\Delta_{\nu}a_{\mu}-2\pi s_{\mu\nu})^{2}-iN\sum_{\ell}n_{\mu}\,a_{\mu}+\frac{1}{2g^{2}}\sum_{\tilde{\ell}}(\Delta_{\mu}\varphi-2\pi s_{\mu})^{2}\\ &\hskip 14.22636pt-iN\sum_{R}n(R)\varphi(R)+\frac{iN\Theta}{8\pi^{2}}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\,(\Delta_{\mu}\varphi-2\pi s_{\mu})(R)(\Delta_{\nu}a_{\lambda}-\Delta_{\lambda}a_{\nu}-2\pi s_{\nu\lambda})(r),\end{split}$ (2.1)
where $\Delta_{\mu}\varphi(R)\equiv\varphi(R+e_{\mu})-\varphi(R)$ denotes the lattice difference along the link between site $R$ and site $R+e_{\mu}$. The gauge field $a_{\mu}\in\mathbb{R}$ and electric particle worldlines $n_{\mu}\in\mathbb{Z}$ are on links $\ell$ between sites $r$, and the Dirac strings $s_{\mu\nu}\in\mathbb{Z}$ of the monopoles live on plaquettes $P$. Meanwhile, $\varphi\in\mathbb{R}$ and $n\in\mathbb{Z}$ live on sites $R$ of the dual lattice, and the $s_{\mu}\in\mathbb{Z}$ are on links $\tilde{\ell}$ of the dual lattice. The last term, a (2+1)d analogue of the theta term in (3+1)d, is an interaction between each plaquette and the link on the dual lattice that intersects it (see Figure 1).
In the way we have expressed Eq. (2.1), we have embedded the $\mathbb{Z}_{N}$ gauge theory and the $\mathbb{Z}_{N}$ matter theory into $U(1)$ theories. Indeed, if $n_{\mu}=n=0$, then the action reduces to that of $U(1)$ gauge theory coupled to an $XY$ model by a topological term. But the gauge group and matter fields should still be regarded as $\mathbb{Z}_{N}$ fields. By the Poisson summation formula, summing over $n_{\mu}$ and $n$ introduce constraints that $\varphi,a_{\mu}\in\frac{2\pi}{N}\mathbb{Z}$. Hence, we can either interpret $n_{\mu}$ and $n$ as electric worldlines and spin operators respectively, or we can view them as Lagrange multipliers that constrain $a_{\mu}$ and $\varphi$ to be $\mathbb{Z}_{N}$ fields.
The lattice model is compact, so there are additionally two classes of disorder operators (for a review see Ref. [52]) in this theory. The vortex currents of the spin model are represented by
$m_{\mu}(r)=\varepsilon_{\mu\nu\lambda}\,\Delta_{\nu}s_{\lambda}$ (2.2)
and are on links of the direct lattice. The monopoles of the gauge theory,
$m(R)=-\frac{1}{2}\varepsilon_{\mu\nu\lambda}\,\Delta_{\mu}s_{\nu\lambda},$ (2.3)
are on sites of the dual lattice. To elucidate the physical role of the topological term, it is useful to express this term as
$S_{\mathrm{topo}}$ $=\frac{iN\Theta}{8\pi^{2}}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\,(\Delta_{\mu}\varphi-2\pi s_{\mu})(\Delta_{\nu}a_{\lambda}-\Delta_{\lambda}a_{\nu}-2\pi s_{\nu\lambda})$ (2.4)
$=\sum_{r,\,R}\left(\frac{iN\Theta}{4\pi^{2}}\varepsilon_{\mu\nu\lambda}\,\Delta_{\mu}\varphi\,\Delta_{\nu}a_{\lambda}+\frac{iN\Theta}{2}\varepsilon_{\mu\nu\lambda}\,s_{\mu}\,s_{\nu\lambda}-\frac{iN\Theta}{2\pi}(m_{\mu}\,a_{\mu}+m\,\varphi)\right).$ (2.5)
The last two terms couple the vortices and monopoles to $a_{\mu}$ and $\varphi$ just like $n_{\mu}$ and $n$ respectively. Thus, the topological term effectively shifts the quantum numbers to
$\begin{split}(n,m)\mapsto\left(n+\frac{\Theta}{2\pi}\,m,m\right),\qquad(n_{\mu},m_{\mu})\mapsto\left(n_{\mu}+\frac{\Theta}{2\pi}\,m_{\mu},m_{\mu}\right).\end{split}$ (2.6)
The vortices of the spin model gain an electric charge proportional to $\Theta$, and the monopoles of the gauge theory obtain a fractional $\mathbb{Z}_{N}$ spin. This phenomenon is the analogue of the Witten effect in (3+1)d, under which a magnetic monopole gains an electric charge proportional to $\Theta$ when there is a theta term [36]. For this reason, we refer to Eq. (2.6) as the generalized Witten effect.
2.2 Global symmetries
As mentioned previously, our model is equipped with two global symmetries, which we now discuss more formally. There is an ordinary (i.e., zero-form) $\mathbb{Z}_{N}$ global symmetry under which $\varphi$ transforms as
$\varphi\to\varphi+\frac{2\pi}{N}.$ (2.7)
The local operator that carries the $\mathbb{Z}_{N}$ charge is
$V(R)=e^{i\varphi(R)}$ (2.8)
on a dual lattice site $R$. Under the symmetry transformation, Eq. (2.7), this operator transforms as
$V(R)\to\omega\,V(R),$ (2.9)
where $\omega$ is an $N$th root of unity. There is also a $\mathbb{Z}_{N}$ one-form symmetry that transforms the gauge field as
$a_{\mu}\to a_{\mu}+\frac{2\pi\eta_{\mu}}{N},$ (2.10)
where $\eta_{\mu}\in\mathbb{Z}$ and $\Delta_{\mu}\eta_{\nu}-\Delta_{\nu}\eta_{\mu}=0$ (i.e., $\eta_{\mu}$ is a flat connection). The gauge invariant operator that transforms under this global symmetry is the Wilson loop,
$W(\Gamma)=\prod_{\ell\,\in\,\Gamma}e^{ia_{\mu}(\ell)},$ (2.11)
which is a product over all links $\ell$ which define the noncontractible loop $\Gamma$. The one-form global symmetry, Eq. (2.10), transforms the Wilson loop to
$W(\Gamma)\to\omega\,W(\Gamma),$ (2.12)
where $\omega$ is again an $N$th root of unity. As mentioned in Section 1, we will use the notation $\mathbb{Z}_{N}^{(q)}$ to denote a $\mathbb{Z}_{N}$ $q$-form symmetry, so our lattice model in Eq. (2.1) then has the global symmetry $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$.
To probe these global symmetries, it is useful to couple to background fields. The $\mathbb{Z}_{N}^{(0)}$ global symmetry couples to a background field $2\pi A_{\mu}/N$ where $A_{\mu}\in\mathbb{Z}$. Similarly, the $\mathbb{Z}_{N}^{(1)}$ couples to a two-form field $2\pi B_{\mu\nu}/N$, where $B_{\mu\nu}\in\mathbb{Z}$ is antisymmetric in its indices. In the presence of background fields, the partition function becomes
$\begin{split}Z[A_{\mu},B_{\mu\nu}]&=\int[da_{\mu}][d\varphi]\sum_{\left\{n,\,n_{\mu},\,s_{\mu},\,s_{\mu\nu}\right\}}\delta(\Delta_{\mu}n_{\mu})\,e^{-S[\varphi,\,a_{\mu},\,n,\,n_{\mu},\,s_{\mu},\,s_{\mu\nu};\,A_{\mu},\,B_{\mu\nu}]},\\ S&=\frac{1}{4e^{2}}\sum_{P}\left(f_{\mu\nu}[B]\right)^{2}-iN\sum_{\ell}n_{\mu}\,a_{\mu}+\frac{1}{2g^{2}}\sum_{\tilde{\ell}}\left(\omega_{\mu}[A]\right)^{2}-iN\sum_{R}n(R)\varphi(R)\\ &\hskip 14.22636pt+\frac{iN\Theta}{8\pi^{2}}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\left(\omega_{\mu}[A]\right)\left(f_{\nu\lambda}[B]\right),\end{split}$ (2.13)
where we have defined
$\begin{split}\omega_{\mu}[A]=\Delta_{\mu}\varphi-2\pi s_{\mu}-\frac{2\pi A_{\mu}}{N},\qquad f_{\mu\nu}[B]=\Delta_{\mu}a_{\nu}-\Delta_{\nu}a_{\mu}-2\pi s_{\mu\nu}-\frac{2\pi B_{\mu\nu}}{N}.\end{split}$ (2.14)
The gauge transformations for the background fields act as
$A_{\mu}\to A_{\mu}+\Delta_{\mu}\chi+N\mathcal{N}_{\mu},\qquad\varphi\to\varphi+\frac{2\pi}{N}\xi,\qquad s_{\mu}\to s_{\mu}-\mathcal{N}_{\mu},$ (2.15)
$B_{\mu\nu}\to B_{\mu\nu}+\Delta_{\mu}\xi_{\nu}-\Delta_{\nu}\xi_{\mu}+N\mathcal{M}_{\mu\nu},\qquad a_{\mu}\to a_{\mu}+\frac{2\pi\xi_{\mu}}{N},\qquad s_{\mu\nu}\to s_{\mu\nu}-\mathcal{M}_{\mu\nu},$
where $\chi,\xi_{\mu},\mathcal{N}_{\mu},\mathcal{M}_{\mu\nu}\in\mathbb{Z}$. From the background fields, $A_{\mu}$ and $B_{\mu\nu}$, we can define corresponding field strengths,
$F_{\mu\nu}=\Delta_{\mu}A_{\nu}-\Delta_{\nu}A_{\mu},\qquad H_{\mu\nu\lambda}=\Delta_{\mu}B_{\nu\lambda}+\Delta_{\nu}B_{\lambda\mu}+\Delta_{\lambda}B_{\mu\nu},$ (2.16)
which are antisymmetric in all their indices. On the lattice, these field strengths can take any value (mod $N$). In the continuum limit, as reviewed in Appendix A, any gauge field for a discrete symmetry must be locally flat (i.e., $F_{\mu\nu}=H_{\nu\lambda\sigma}=0$ mod $N$). Since our interest is ultimately in the continuum limit, we will require $A_{\mu}$ and $B_{\mu\nu}$ to be locally flat. In the next section, we will see that coupling to probes for the global symmetries is an important tool that will help us determine the phases and phase diagram.
2.3 Duality
Before turning to the phase diagram of our lattice model, Eq. (2.1), it is useful to determine any non-perturbative information we can exploit to constrain the phase diagram. The model turns out to have two kinds of duality transformations. First, we note that in the absence of background fields, the partition function is periodic under shifts of $\Theta$ by integer multiples of $2\pi$. If we also introduce background gauge fields and perform the shift $\Theta\to\Theta+2\pi$, the action in Eq. (2.13) changes by
$\Delta S$ $=\frac{iN2\pi}{8\pi^{2}}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\left(\Delta_{\mu}\varphi-2\pi s_{\mu}-\frac{2\pi A_{\mu}}{N}\right)\left(\Delta_{\nu}a_{\lambda}-\Delta_{\lambda}a_{\nu}-2\pi s_{\nu\lambda}-\frac{2\pi B_{\nu\lambda}}{N}\right)$ (2.17)
$=\frac{iN}{2\pi}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\,\Delta_{\mu}\varphi\,\Delta_{\nu}a_{\lambda}+i2\pi N\sum_{r,\,R}\frac{1}{2}\varepsilon_{\mu\nu\lambda}\left[s_{\mu}\left(s_{\nu\lambda}+\frac{B_{\nu\lambda}}{N}\right)+\varepsilon_{\mu\nu\lambda}\frac{A_{\mu}}{N}s_{\nu\lambda}\right]$ (2.18)
$\hskip 14.22636pt-iN\sum_{R}\left(m-\frac{1}{2N}\,\varepsilon_{\mu\nu\lambda}\,\Delta_{\mu}B_{\nu\lambda}\right)\varphi-iN\sum_{\ell}\left(m_{\mu}+\frac{1}{N}\,\varepsilon_{\mu\nu\lambda}\,\Delta_{\nu}A_{\lambda}\right)a_{\mu}$
$\hskip 14.22636pt+\frac{iN}{2\pi}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\frac{2\pi A_{\mu}}{N}\frac{1}{2}\frac{2\pi B_{\nu\lambda}}{N}.$
The first term in the first line of Eq. (2.18) is a total lattice derivative and thus will not contribute if we take periodic boundary conditions. The remaining terms in the first line of Eq. (2.18) are integer multiples of $2\pi i$, so they also can be ignored. The terms in the second line of Eq. (2.18) may be removed by correspondingly shifting
$n\to n-m+\frac{1}{2N}\,\varepsilon_{\mu\nu\lambda}\,\Delta_{\mu}B_{\nu\lambda},\qquad n_{\mu}\to n_{\mu}-m_{\mu}-\frac{1}{N}\,\varepsilon_{\mu\nu\lambda}\,\Delta_{\nu}A_{\lambda}.$ (2.19)
Because the background fields are locally flat, Eq. (2.19) is a shift of $n$ and $n_{\mu}$ by integers, so such a transformation is well-defined. After the shift $\Theta\mapsto\Theta+2\pi$, the only term we cannot remove is
$\Delta S=\frac{iN}{2\pi}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\frac{2\pi A_{\mu}}{N}\frac{1}{2}\frac{2\pi B_{\nu\lambda}}{N}.$ (2.20)
Thus, although the partition function is invariant under shifting $\Theta$ by $2\pi$ in the absence of background fields, it gains an additional phase factor in the presence of background fields. Hence, we may regard shifting $\Theta$ by $2\pi$ as stacking a $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$ SPT.
Next, we turn to the other duality transformation, which is analogous to Kramers-Wannier duality. As demonstrated in Appendix C, by manipulating the partition function, Eq. (2.13), in ways similar to Refs. [7, 53], we can find a dual theory with action,
$\begin{split}S&=\frac{1}{4\tilde{e}^{2}}\sum_{P}(\tilde{f}_{\mu\nu}[b])^{2}-iN\sum_{\ell}\tilde{n}_{\mu}\,\tilde{a}_{\mu}+\frac{1}{2\tilde{g}^{2}}\sum_{\tilde{\ell}}\left(\tilde{\omega}_{\mu}[c]\right)^{2}-iN\sum_{R}\tilde{n}(R)\tilde{\varphi}(R)\\ &\hskip 14.22636pt-\frac{iN\widetilde{\Theta}}{8\pi^{2}}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\,\tilde{\omega}_{\mu}[c]\,\tilde{f}_{\nu\lambda}[b]+\frac{\pi i}{N}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\left(b_{\nu\lambda}\,\Delta_{\mu}\tilde{k}+k_{\mu}\,\Delta_{\nu}c_{\lambda}\right)\\ &\hskip 14.22636pt+\frac{\pi i}{N}\sum_{r,\,R}\varepsilon_{\mu\nu\lambda}\,\left(c_{\mu}\,B_{\nu\lambda}+b_{\mu\nu}\,A_{\lambda}\right),\end{split}$ (2.21)
where we use the abbreviated notation,
$\tilde{\omega}_{\mu}[c]=\Delta_{\mu}\tilde{\varphi}-2\pi\tilde{s}_{\mu}-\frac{2\pi c_{\mu}}{N},\qquad\tilde{f}_{\mu\nu}[b]=\Delta_{\mu}\tilde{a}_{\nu}-\Delta_{\nu}\tilde{a}_{\mu}-2\pi\tilde{s}_{\mu\nu}-\frac{2\pi b_{\mu\nu}}{N},$ (2.22)
and we have defined
$\frac{\tilde{e}^{2}}{e^{2}}=\frac{\tilde{g}^{2}}{g^{2}}=\left(\frac{\Theta}{2\pi}\right)^{2}+\left(\frac{2\pi}{Nge}\right)^{2},\qquad\widetilde{\Theta}=-\frac{\Theta}{\left(\frac{\Theta}{2\pi}\right)^{2}+\left(\frac{2\pi}{Nge}\right)^{2}}.$ (2.23)
The gauge field $\tilde{a}_{\mu}\in\mathbb{R}$ is on links of the direct lattice, and $\tilde{\varphi}\in\mathbb{R}$ is on sites of the dual lattice. We also have $\tilde{n}_{\mu}\in\mathbb{Z}$ lives on links and $\tilde{n}\in\mathbb{Z}$ on dual sites. The variables $\tilde{s}_{\mu\nu},b_{\mu\nu}\in\mathbb{Z}$ are on plaquettes while $\tilde{s}_{\mu},c_{\mu}\in\mathbb{Z}$ are on dual links. Finally, $k_{\mu}\in\mathbb{Z}$ are on links, and $\tilde{k}\in\mathbb{Z}$ are on dual sites. All these lattice variables are dynamical except $A_{\mu}$ and $B_{\mu\nu}$, which are probes.
The duality of Eq. (2.21) and Eq. (2.13) demonstrates that our lattice model is locally self-dual. Indeed, Eq. (2.21) is similar to Eq. (2.13) but has two important differences. First, the coupling constants have been modified. This mapping of coupling constants is nicely packaged if we define the complex coupling constant $\tau$ and the coupling ratio $\kappa$ to be
$\tau=\frac{\Theta}{2\pi}+i\frac{2\pi}{Nge},\qquad\kappa=\frac{e}{g}.$ (2.24)
Duality then maps the coupling constants as
$\mathcal{S}:\qquad\tau\to-\frac{1}{\tau},\qquad\kappa\to\kappa.$ (2.25)
Secondly, a more subtle difference is that the flat background gauge fields in Eq. (2.13) are now dynamical in Eq. (2.21), so the lattice model is dual to a gauged version of itself. This duality is a generalization of the duality relating a $\mathbb{Z}_{N}$ clock model and a $\mathbb{Z}_{N}$ gauge theory in (2+1)d.
To summarize, we have two kinds of duality transformations. The partition function $Z[\tau;A_{\mu},B_{\mu\nu}]$ depends on [^4] The partition function also depends on $\kappa=e/g$, but we suppress that dependence since $\kappa$ is invariant under $\mathcal{T}$ and $\mathcal{S}$ in Eq. (2.26). the complex coupling constant $\tau$ and the background fields $A_{\mu}$ and $B_{\mu\nu}$. We then define transformations $\mathcal{T}$ and $\mathcal{S}$ of the partition function as
$\begin{split}\mathcal{T}\left(Z[\tau;A_{\mu},B_{\mu\nu}]\right)&=Z[\tau+1;A_{\mu},B_{\mu\nu}]=Z[\tau;A_{\mu},B_{\mu\nu}]e^{-\frac{\pi i}{N}\sum_{r,R}\varepsilon_{\mu\nu\lambda}\,A_{\mu}\,B_{\nu\lambda}},\\ \mathcal{S}\left(Z[\tau;A_{\mu},B_{\mu\nu}]\right)&=Z\left[-1/\tau;A_{\mu},B_{\mu\nu}\right]=\sum_{\left\{a_{\mu},\,b_{\mu\nu}\right\}}Z[\tau;a_{\mu},b_{\mu\nu}]e^{-\frac{\pi i}{N}\sum_{r,R}\varepsilon_{\mu\nu\lambda}\left(a_{\mu}\,B_{\nu\lambda}+b_{\mu\nu}\,A_{\lambda}\right)}.\end{split}$ (2.26)
where $a_{\mu},b_{\mu\nu}\in\mathbb{Z}$ are locally flat dynamical $\mathbb{Z}_{N}$ gauge fields while $A_{\mu},B_{\mu\nu}\in\mathbb{Z}$ are locally flat background $\mathbb{Z}_{N}$ gauge fields. It is also useful to consider the continuum analogues of the transformations $\mathcal{S}$ and $\mathcal{T}$. The analogous transformations for a Euclidean action $S$ are
$\begin{split}\mathcal{T}:&\qquad S[A,B]\to S[A,B]+\frac{iN}{2\pi}\int A\wedge B,\\ \mathcal{S}:&\qquad S[A,B]\to S[a,b]+\frac{iN}{2\pi}\int(a\wedge B+b\wedge A),\end{split}$ (2.27)
where $a_{\mu}$ and $b_{\mu\nu}$ are dynamical $\mathbb{Z}_{N}$ gauge fields while $A_{\mu}$ and $B_{\mu\nu}$ are background $\mathbb{Z}_{N}$ gauge fields. We will use both Eqs. (2.26) and (2.27) interchangeably. The mathematical structure of the $\mathcal{S}$ and $\mathcal{T}$ transformations is reminiscent of that in the fractional quantum Hall effect and (2+1)d conformal field theories [54, 55]. A similar structure also arises in (3+1)d lattice gauge theory with a theta term [41].
As we demonstrate in Appendix C, we can deduce how the $\mathcal{S}$ and $\mathcal{T}$ transformations map between different operators, allowing us determine what objects condense in a given phase. The local operators are labeled $(q_{s},q_{m})$ by a spin charge $q_{s}$ and a monopole charge $q_{m}$, and the loop operators are labeled $(q_{e},q_{v})$ by an electric charge $q_{e}$ and a vorticity $q_{v}$. Because the dynamical spins and electric charges are multiples of $N$, we note that a local operator that can condense must have $q_{s}\in N\mathbb{Z}$, and a loop operator that condense must have an electric charge $q_{e}\in N\mathbb{Z}$. The $\mathcal{S}$ and $\mathcal{T}$ transformations act on these objects as
[…]
We can then place many constraints on the phase diagram using these transformations. For example, if a certain phase has local operators with $(q_{s},q_{m})$ condensed, then the image of this phase under $\mathcal{S}$ will have local operators with charges $(-Nq_{m},q_{s}/N)$ condensed. The phase diagram can then be deduced by acting with a series of $\mathcal{S}$ and $\mathcal{T}$ transformations on the phases near $\Theta=0$.
2.4 Phases and phase diagram
Figure 2: The phase diagram at $\Theta=0$ for fixed $\kappa=e/g$ as $\mathrm{Im}(\tau)=2\pi/Nge$ is varied. In this limit, the lattice model reduces to a decoupled $\mathbb{Z}_{N}$ clock model and $\mathbb{Z}_{N}$ gauge theory. At large $e$ and $g$, there is a trivial phase in which the clock model is disordered and the gauge theory is confining, fully preserving the global symmetry $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$. At small $e$ and $g$, the clock model orders, and the gauge theory is deconfined (and topologically ordered), breaking the full symmetry group $G$. Depending on the value of $\kappa=e/g$, there may be an intermediate phase in which one of $\mathbb{Z}_{N}^{(0)}$ or $\mathbb{Z}_{N}^{(1)}$ is broken but not the other.
We are now equipped to understand the phases of the model and how the phase diagram is organized. We begin with the phase structure at $\Theta=0$, where we have a decoupled $\mathbb{Z}_{N}$ clock model and $\mathbb{Z}_{N}$ gauge theory. Here, the physics is well understood. The clock model has two phases—the ordered phase and the disordered phase. At small $g$, the $\mathbb{Z}_{N}$ spins condense into an ordered phase in which the N(0